OpenStudy (anonymous):

Solve for x: 4 − (x + 2) < −3(x + 4) x < −7 x > −7 x < −9 x > −9

OpenStudy (imstuck):

That goes like this:\[4-x-2<-3x-12\]then\[2-x <-3x-12\]

OpenStudy (anonymous):

Can you explain each step? This is from a test and I'm really stuck.

OpenStudy (imstuck):

\[14<-2x\]

OpenStudy (imstuck):

Because you are dividing by a negative number, you have to change the sign when you do the division?

OpenStudy (imstuck):

Step by step? Ok, here goes...

OpenStudy (anonymous):

please

OpenStudy (imstuck):

Ok starting right at the start with the original inequality:

OpenStudy (imstuck):

\[4-(x+2)<-3(x+4)\]The negative sign on the left changes the sign inside the parenthesis. Like this:\[4-x-2<-3(x+4)\]

OpenStudy (imstuck):

Simplifying the left side, 4 - 2 = 2 so \[2-x <-3(x+4)\]

OpenStudy (imstuck):

The right side then needs to distribute the -3 into the parenthesis:\[2-x <-3x-12\]

OpenStudy (imstuck):

Now you combine the x's on one side and the numbers on the other, like this:\[2+12<-3x+x\]

OpenStudy (imstuck):

Simplifying gives you this:\[14<-2x\]

OpenStudy (anonymous):

is it x > 7

OpenStudy (imstuck):

Like I said before, when you divide by a negative number you have to change the inequality sign:\[-\frac{ 14 }{ 2 }>\frac{ -2 }{ -2 }x\]

OpenStudy (imstuck):

\[-7>x\]which is the same thing as x<-7. Your first choice from your list of answers.

OpenStudy (imstuck):

And that's it! You're done!

OpenStudy (anonymous):

Sorry i never gave you a medal, I kind of passed out last night. theweed2stronk @IMstuck